# Enhancing Training Data Attribution with Representational Optimization

### v1

@@ Abstract
Training data attribution (TDA) methods aim to measure how training data impacts a model’s predictions. While gradient-based attribution methods, such as influence functions, offer theoretical grounding, their computational costs make them impractical for large-scale applications. Representation-based approaches are far more scalable, but typically rely on heuristic embeddings that are not optimized for attribution, limiting their fidelity. To address these challenges, we propose AirRep, a scalable, representation-based approach that closes this gap by learning task-specific and model-aligned representations optimized explicitly for TDA. AirRep introduces two key innovations: a trainable encoder tuned for attribution quality, and an attention-based pooling mechanism that enables accurate estimation of group-wise influence. We train AirRep using a ranking objective over automatically constructed training subsets labeled by their empirical effect on target predictions. Experiments on instruction-tuned LLMs demonstrate that AirRep achieves performance on par with state-of-the-art gradient-based approaches while being nearly two orders of magnitude more efficient at inference time. Further analysis highlights its robustness and generalization across tasks and models. Our code is available at github.com/sunnweiwei/AirRep.
@@ 1 Introduction
The remarkable success of large language models (LLMs) has been demonstrated across a wide range of tasks. However, a fundamental question remains open in machine learning: how does the behavior of LLMs depend on their training data? More specifically, what training examples cause models to generalize well—or underperform—on specific inputs or tasks? Training Data Attribution (TDA), the process of measuring the impact of specific parts of the training data on the predictions of machine learning models, is an important step to answering this question [1]. TDA is crucial for ensuring transparency and accountability in AI systems by revealing how data influences model outputs.
Existing approaches to TDA can broadly be categorized as either gradient-based or representation-based. Gradient-based approaches, rooted in influence estimation [2, 3], aim to quantify the impact of individual training points on model predictions. Their core idea is to use gradients and the inverse Hessian of the loss function to make a first-order approximation of how a model’s predictions would change if a certain example was removed from the training set [3]. While these methods are theoretically well-motivated, they are typically computationally expensive and rely on assumptions of loss convexity and model optimality — both of which are violated in modern neural networks [4, 5].
In contrast, representation-based methods estimate influence based on similarity in the representation space. The core assumption is that training examples whose representations are close to a test input are likely to have influenced its prediction [6]. Recent works have explored various notions of similarity based on model hidden states [7, 6], n-gram features [8], and text embeddings [9, 10]. Compared to gradient-based TDA, representation-based approaches are more computationally efficient and scalable, making them well-suited for large-scale applications, such as curating pre-training data for LLMs [8, 11]. However, the quality of a representation-based method depends heavily on the selected feature space. This problem is compounded by the fact that existing work utilizes heuristically designed representations that are not tailored to the particular target task or model [12].
Finally, a persistent challenge in both method families is is how these methods are extended to estimate the collective influence of a group of training examples. Most past work extends single-example attribution methods to the group setting by simply summing over individual attribution scores within the group. However, this additive assumption fails to capture groupwise interactions and may lead to inaccurate influence estimates [13].
@@
Figure 1: Performance comparison of gradient-based and representation-based training data attribution (TDA) approaches. Left: Average linear data model score (LDS) [1] on 5 unseen datasets (FLAN, Alpaca, Tulu, SafeRLHF). AirRep outperforms state-of-the-art gradient-based methods such as LoGra (which uses 48$\times$ more storage). Right: Inference speed (encoded examples per second on a single GPU). AirRep is nearly two orders of magnitude more efficient than gradient-based methods during inference.
@@ 2 Preliminaries
where $\ell(\cdot)$ denotes the cross-entropy token prediction loss. As language model training is non-convex, we use $\theta^{*}$ to denote the optimized parameters after fine-tuning, regardless of whether a true minimum has been found. We define the model’s “prediction” on $x$ as the cross-entropy loss computed on $x$, given by $r(x,S)=\ell(x;\theta^{*})$, i.e. the evaluation loss on the test example $x$ for the model fine-tuned on the dataset $S$. Here, we use the cross-entropy loss instead of task-specific metrics like the accuracy, F1 score, or ROUGE [18] because the loss is a generic metric that is defined for all NLP tasks and has been found to be correlated with task-specific metrics [19].
A data attribution model, denoted by $f(x,S)$, aims to estimates the actual re-training outcome $r(x,S)$. Existing methods can be generally categorized into two groups: gradient-based methods and representation-based methods.
@@
Gradient-based methods largely stem from influence functions [2, 3], which approximate how $\theta$ would change in response to infinitesimal perturbations in the weighting of training instances. Specifically, influence functions quantify the influence from a single training example $z_{i}$ to an evaluation example $x$ with a closed form expression:
$f_{\text{IF}}(x,z_{i})=-\nabla_{\theta}\ell(x;\theta)\;\mathbf{H}^{-1}\;\nabla_{\theta}\ell(z_{i};\theta),$ (2)
where $\mathbf{H}^{-1}$ is the inverse Hessian of the training loss wrt model parameters. The derivation of Eq. 2 is given by first-order Taylor approximation of $\ell(x;\theta)$ around $\theta^{*}$ [20] (see Appendix A for details).
The size of model neural networks makes the computation and inversion of the Hessian intractable. Consequently, practical influence function-based methods often use an approximation of the Hessian [12, 21]. Recent techniques have also considered techniques like gradient projection [16] and model ensembling [22, 23]. As a representative approach in LLM application, LoGra [16] define the influence of a training example $z_{i}$ on $x$ is computed as:
@@
where $\operatorname{norm}\bigl{[}\mathbf{z}\bigl{]}_{2}=\frac{\mathbf{z}}{\|\mathbf{z}\|_{2}}$ is a unit normalization operation [24]; $\hat{\theta}$ denotes the trainable weights of a LoGra module [16] with PCA initialization for gradient projection; $\nabla_{\hat{\theta}}\ell(z;\theta)$ is the gradient of the loss with respect to $\hat{\theta}$; and $\mathbf{H_{\hat{\theta}}}^{-\frac{1}{2}}$ is computed on $\hat{\theta}$ and is approximated based on the Kronecker-Factored Approximate Curvature (KFAC) algorithm [21]. Further, under first-order approximation, the group influence $f_{\text{GD}}(x,S)$ is estimated as the summation of individual influences.
@@
where $\operatorname{Enc}(z)$ denotes a continuous embedding of input $z$ encoded by encoder model $\operatorname{Enc}$. Group influence is usually defined as summation of individual influence under a linear assumption [1].
In term of the encoding function $\operatorname{Enc}$, existing work in the text domain has investigated different heuristic designs including (last-layer) hidden states of a language model [7, 6], n-gram representations [8], TF-IDF representations [25, 26], and text embeddings [11, 27]. See Appendix C for further comparison of the two types of TDA methods.
@@ 3 AirRep
Specifically, AirRep incorporates two novel mechanisms to improve performance over standard representation-based methods: First, we introduce an effective attention-based pooling mechanism that better reflects group influence effect (Section 3.1). Second, we optimize the model on auto-generated data so that AirRep’s score better reflects the underlying models predictions (Section 3.2).
@@ 3.1 Attention-based Pooling
$\text{where}\quad\alpha_{i}$ $=\frac{\exp(|\operatorname{Enc}(x)^{\top}\cdot\operatorname{Enc}(z_{i})|)}{\sum_{j\in[n]}\exp(|\operatorname{Enc}(x)^{\top}\cdot\operatorname{Enc}(z_{i})|)}.$
where $\operatorname{Enc}(x)$ denotes the sentence embedding predicted by a BERT-based sentence encoder [28, 29], $\alpha_{i}$ denotes the attention score of example $z_{i}$, and ||$\cdot$|| ⋅ | denotes the absolute value operation.
@@
Empirically, we observe that influence scores are sparse, where each test example relies on only a few training points, while others add noise. This support the need for selective pooling, consistent with prior findings [1, 12, 30]. Additionally, the proposed attention-based influence pooling can also related to high-order group influence functions. Specifically, Basu et al. [13] showed that second-order terms capture additional relationship between samples. Extending this, we show that high-order group influence introduces sample-wise weights, akin to our attention-based pooling (see Appendix B.1 for more discussion).
@@ 3.2 Optimization
In practice, the label $r_{i}$ is noisy due to the stochastic nature of LM training. To mitigate the impact of noisy labels, we adopt a weighted pairwise ranking loss, inspired by importance reweighting [31, 32], which assigns lower weights to uncertain labels while prioritizing more reliable ones:
$\small\begin{split}\mathcal{L}(x,\mathcal{S})&=-\sum_{i,j\in M}\mathds{1}_{r_{i}>r_{j}}\,w_{i,j}\,\log\sigma(f_{i}-f_{j}),\\ \text{where}&\quad w_{i,j}=\begin{cases}0,&\text{if }|r_{i}-r_{j}|<T_{\min},\\ \min\{|r_{i}-r_{j}|,T_{\max}\},&\text{if }T_{\min}\leq|r_{i}-r_{j}|.\end{cases}\end{split}$ (9)
@@
TracIn [22] 18432 (48$\times$) 11.33 14.75 9.21 10.75 10.60
LESS [25] 8196 (21$\times$) 16.16 16.40 9.59 13.02 25.63
LoGra [16] 1152   (3$\times$) 13.78 13.32 6.87 10.16 24.76
LoGra [16] 18432 (48$\times$) 18.45 19.75 12.38 14.88 26.82
Dsdm [33] 18432 (48$\times$) 18.02 19.67 12.15 14.31 25.94
TF-IDF [34] - 9.98 2.52 7.24 5.24 24.94
DSIR [8] - -0.02 0.49 2.01 -0.49 -2.10
RDS [6] 896 (2.3$\times$) 3.86 0.74 0.87 1.89 11.94
GTE-Small [35] 384    (1$\times$) 7.65 0.92 1.74 1.14 26.80
AirRep (Ours) 384    (1$\times$) 26.23 21.11 22.58 15.14 46.08
@@ 4 Experimental Setup
Our experiments focus on LM finetuning, using the Qwen2.5 model family [36] as our base LMs. During training, we start with the base LM and fine-tune it using a batch size of $32$ and the AdamW optimizer [37] with a learning rate of 2e-5 for two epochs.
@@
For AirRep training, we utilize two datasets, each representing distinct scenarios: instruction tuning on standard NLP tasks and training to improve conversational abilities: (i) FLAN [14], an instruction-tuning dataset for language models, and (ii) UltraChat [15], a large-scale dataset of instructional conversations covering a wide range of topics. To generate training signal, we set $N_{v}=10^{4}$ and $N_{t}=10^{5}$. The training subsets number is $M=100$, with each subsets containing $n=1{,}000$ samples. We construct $100$ cross-validation instances. Thus, in total, the data includes $10^{4}$ unique training subsets and $10^{7}$ training examples. The Qwen2.5-0.5B LM is then fine-tuned on these training subsets and evaluated on the corresponding validation subset to obtain the label $r(\cdot,\cdot)$.
@@
We initialize AirRep using the GTE-Small, a 30M parameter embedding model [35], and apply a randomly initialized projection matrix on top. AirRep is trained separately on FLAN and UltraChat. To construct the data for each training step of AirRep, we randomly select one cross-validation instance, then sample $1{,}000$ examples from its validation subset and $32$ training subsets. Distributed training is employed to maximize GPU memory utilization. The clipping thresholds, $T_{\text{min}}$ and $T_{\text{max}}$, are set to $0.1$ and $5.0$, respectively. The model is optimized for up to $2{,}000$ steps using the AdamW optimizer with a learning rate of $1\times 10^{-4}$.
@@
We use the following datasets for evaluation: (i) FLAN [14]: The evaluation data of FLAN contains 66 NLP tasks spanning diverse categories. (ii) Alpaca [38]: An instruction-tuning dataset generated by OpenAI’s text-davinci-003 model. (iii) Tulu [39]: An instruction-tuning dataset comprising diverse data sources. (iv) SafeRLHF [40]: A dataset for safety alignment of large language models, where each response is labeled as either safe or unsafe. Note that each evaluation data contain a test set and a training set. We ensure that all evaluation data remain excluded from AirRep’s optimization data to guarantee that the evaluation results reflect AirRep’s generalization to unseen data. Appendix D provides additional dataset details.
@@
We compare AirRep with representative gradient-based and representation-based methods. The gradient-based baselines we consider are as follows: (i) LoGra [16], an optimized version of the influence function as described in Eq.4, based on its implementation in the Logix software library. (ii) TracIn [22], which leverages the dot product of gradient vectors. (iii) LESS [25], which computes projected gradients on LoRA weights and adjusts gradients using the AdamW states. (iv) Dsdm [33], an efficient implementation of influence function with FIM.
The representation-based baselines we consider are as follows: (i) TF-IDF [34], (ii) DSIR [8], hashed n-gram features, (iii) RDS [6], which utilizes the last-layer hidden states of LMs, and (iv) GTE [35], a text embedding model trained on text similarity and retrieval tasks which was also used as the base pre-trained model for AirRep. See Appendix E for more implementation details.

### v2

@@ Abstract
Training data attribution (TDA) methods aim to measure how training data impacts a model’s predictions. While gradient-based attribution methods, such as influence functions, offer theoretical grounding, their computational costs make them impractical for large-scale applications. Representation-based approaches are far more scalable, but typically rely on heuristic embeddings that are not optimized for attribution, limiting their fidelity. To address these challenges, we propose AirRep, a scalable, representation-based approach that closes this gap by learning task-specific and model-aligned representations optimized explicitly for TDA. AirRep introduces two key innovations: a trainable encoder tuned for attribution quality, and an attention-based pooling mechanism that enables accurate estimation of group-wise influence. We train AirRep using a ranking objective over automatically constructed training subsets labeled by their empirical effect on target predictions. Experiments on instruction-tuned LLMs demonstrate that AirRep achieves performance on par with state-of-the-art gradient-based approaches while being nearly two orders of magnitude more efficient at inference time. Further analysis highlights its robustness and generalization across tasks and models. Our code is available at https://github.com/sunnweiwei/AirRep.
@@ 1 Introduction
The remarkable success of large language models (LLMs) has been demonstrated across a wide range of tasks. However, a fundamental question remains open in machine learning: how does the behavior of LLMs depend on their training data? More specifically, what training examples cause models to generalize well—or underperform—on specific inputs or tasks? Training Data Attribution (TDA), the process of measuring the impact of specific parts of the training data on the predictions of machine learning models, is an important step to answering this question [Ilyas2022DatamodelsPP]. TDA is crucial for ensuring transparency and accountability in AI systems by revealing how data influences model outputs.
Existing approaches to TDA can broadly be categorized as either gradient-based or representation-based. Gradient-based approaches, rooted in influence estimation [Hampel1974TheIC, Koh2017UnderstandingBP], aim to quantify the impact of individual training points on model predictions. Their core idea is to use gradients and the inverse Hessian of the loss function to make a first-order approximation of how a model’s predictions would change if a certain example was removed from the training set [Koh2017UnderstandingBP]. While these methods are theoretically well-motivated, they are typically computationally expensive and rely on assumptions of loss convexity and model optimality — both of which are violated in modern neural networks [Basu2020InfluenceFI, Schioppa2023TheoreticalAP].
In contrast, representation-based methods estimate influence based on similarity in the representation space. The core assumption is that training examples whose representations are close to a test input are likely to have influenced its prediction [Hanawa2021EvaluationOS]. Recent works have explored various notions of similarity based on model hidden states [Zhang2018TheUE, Hanawa2021EvaluationOS], n-gram features [Xie2023DataSF], and text embeddings [Akyrek2022TowardsTF, Das2023DEFTUCSDE]. Compared to gradient-based TDA, representation-based approaches are more computationally efficient and scalable, making them well-suited for large-scale applications, such as curating pre-training data for LLMs [Xie2023DataSF, Penedo2024TheFD]. However, the quality of a representation-based method depends heavily on the selected feature space. This problem is compounded by the fact that existing work utilizes heuristically designed representations that are not tailored to the particular target task or model [Park2023TRAKAM].
Finally, a persistent challenge in both method families is is how these methods are extended to estimate the collective influence of a group of training examples. Most past work extends single-example attribution methods to the group setting by simply summing over individual attribution scores within the group. However, this additive assumption fails to capture groupwise interactions and may lead to inaccurate influence estimation [Basu2019OnSG].
@@
Building on the setup of datamodels [Ilyas2022DatamodelsPP], we apply our approach on FLAN [Wei2021FinetunedLM], an instruction-tuning dataset, and UltraChat [Ding2023EnhancingCL], a large-scale dialogue generation dataset, and evaluate the model on five unseen instruction-tuning test set (FLAN, Alapca, Tulu, SafeRLHF) that do not appear in the training data. Figure 1 shows the main evaluation results. Notably, in the standard Linear Datamodeling Score (LDS) evaluation [Ilyas2022DatamodelsPP], AirRep significantly outperforms existing gradient-based approach [Choe2024WhatIY], despite being ${\sim}80\times$ more computationally efficient and ${\sim}50\times$ more storage efficient at inference time. Our evaluation additionally shows consistent patterns across various downstream TDA tasks, including data selection, data source identification, and task classification. AirRep also exhibits strong generalizabilty to new tasks and models, which further underscoring the effectiveness of AirRep (Section 5). Finally, we show that the inference efficiency of AirRep can effectively amortizes the cost of AirRep training [Covert2024StochasticAA] (Section 9).
@@ 2 Preliminaries
where $\ell(\cdot)$ denotes the cross-entropy token prediction loss. As language model training is non-convex, we use $\theta^{*}$ to denote the optimized parameters after fine-tuning, regardless of whether a true minimum has been found. We define the model’s “prediction outcome” on $x$ as the cross-entropy loss computed on $x$, given by $r(x,S)=\ell(x;\theta^{*})$, i.e. the evaluation loss on the test example $x$ for the model fine-tuned on the dataset $S$. Here, we use the cross-entropy loss instead of task-specific metrics like the accuracy, F1 score, or ROUGE [Lin2004ROUGEAP] because the loss is a generic metric that is defined for all NLP tasks and has been found to be correlated with task-specific metrics [Xia2022TrainingTO].
A data attribution model, denoted by $f(x,S)$, aims to estimate the actual retraining outcome r(x,S)r(x,S)—i.e., the loss of the model on example $x$ after training on dataset $S$. Existing methods can be generally categorized into two groups: gradient-based methods and representation-based methods.
@@
Gradient-based methods largely stem from influence functions [Hampel1974TheIC, Koh2017UnderstandingBP], which approximate how $\theta$ would change in response to infinitesimal perturbations in the weighting of training instances. Specifically, influence functions quantify the influence from a single training example $z_{i}$ to an evaluation example $x$ with a closed form expression:
$f_{\text{IF}}(x,z_{i})=-\nabla_{\theta}\ell(x;\theta)^{\top}\,\mathbf{H}^{-1}\,\nabla_{\theta}\ell(z_{i};\theta).$ (2)
where $\mathbf{H}^{-1}$ is the inverse Hessian of the training loss wrt model parameters. The derivation of Eq. 2 is given by first-order Taylor approximation of $\ell(x;\theta)$ around $\theta^{*}$ [Krantz2002TheIF] (see Appendix A for details).
The size of model neural networks makes the computation and inversion of the Hessian intractable. Consequently, practical influence function-based methods often use an approximation of the Hessian [Park2023TRAKAM, Grosse2023StudyingLL]. Recent techniques have also considered techniques like gradient projection [Choe2024WhatIY] and model ensembling [Pruthi2020EstimatingTD, NEURIPS2023_1297ca5c]. As a representative approach in LLM application, LoGra [Choe2024WhatIY] define the influence of a training example $z_{i}$ on $x$ is computed as:
@@
where $\operatorname{norm}\bigl[\mathbf{z}\bigl]_{2}=\frac{\mathbf{z}}{\|\mathbf{z}\|_{2}}$ is a unit normalization operation [Barshan2020RelatIFIE]; $\hat{\theta}$ denotes the trainable weights of a LoGra module [Choe2024WhatIY] with PCA initialization for gradient projection; $\nabla_{\hat{\theta}}\ell(z;\theta)$ is the gradient of the loss with respect to $\hat{\theta}$; and $\mathbf{H_{\hat{\theta}}}^{-\frac{1}{2}}$ is computed on $\hat{\theta}$ and is approximated based on the Kronecker-Factored Approximate Curvature (KFAC) algorithm [Grosse2023StudyingLL]. Further, under first-order approximation, the group influence $f_{\text{GD}}(x,S)$ is estimated as the summation of individual influences.
@@
where $\operatorname{Enc}(z)$ denotes a continuous embedding of input $z$ encoded by encoder model $\operatorname{Enc}$. Group influence is usually defined as summation of individual influence under a linear assumption [Ilyas2022DatamodelsPP].
In term of the encoding function $\operatorname{Enc}$, existing work in the text domain has investigated different heuristic designs including (last-layer) hidden states of a language model [Zhang2018TheUE, Hanawa2021EvaluationOS], n-gram representations [Xie2023DataSF], TF-IDF representations [Xia2024LESSSI, Robertson2009ThePR], and text embeddings [Penedo2024TheFD, Li2024DataCompLMIS]. See Appendix C for further comparison of the two types of TDA methods.
@@ 3 AirRep
Specifically, AirRep incorporates two novel mechanisms to improve performance over standard representation-based methods: First, we introduce an effective attention-based pooling mechanism that better reflects group influence effect (Section˜3.1). Second, we optimize the model on auto-generated data so that AirRep’s score better reflects the underlying models predictions (Section˜3.2).
@@ 3.1 Attention-based Pooling
$\text{where}\qquad\alpha_{i}$ $=\frac{\exp(|\operatorname{Enc}(x)^{\top}\cdot\operatorname{Enc}(z_{i})|)}{\sum_{j\in[n]}\exp(|\operatorname{Enc}(x)^{\top}\cdot\operatorname{Enc}(z_{i})|)}.$
where $\operatorname{Enc}(x)$ denotes the sentence embedding predicted by a BERT-based sentence encoder [Devlin2019BERTPO, Reimers2019SentenceBERTSE], $\alpha_{i}$ denotes the attention score of example $z_{i}$, and $|\cdot|$ denotes the absolute value operation.
@@
Empirically, we observe that influence scores are sparse, where each test example relies on only a few training points, while others add noise. This supports the need for selective pooling, consistent with prior findings [Ilyas2022DatamodelsPP, Park2023TRAKAM, Chang2024ScalableIA]. Additionally, the proposed attention-based influence pooling can also related to high-order group influence functions. Specifically, Basu et al. [Basu2019OnSG] showed that second-order terms capture additional relationship between samples. Extending this, we show that high-order group influence introduces sample-wise weights, akin to our attention-based pooling (see Appendix B.1 for more discussion).
@@ 3.2 Optimization
In practice, the label $r_{i}$ is noisy due to the stochastic nature of LM training. To mitigate the impact of noisy labels, we adopt a weighted pairwise ranking loss, inspired by importance reweighting [liu2015classification, Song2020LearningFN], which assigns lower weights to uncertain labels while prioritizing more reliable ones:
$\small\begin{split}\mathcal{L}(x,\mathcal{S})&=-\sum_{i,j\in M}\mathds{1}_{r_{i}>r_{j}}\,w_{i,j}\,\log\sigma(f_{i}-f_{j}),\\ \text{where}&\qquad w_{i,j}=\begin{cases}0,&\text{if }|r_{i}-r_{j}|<T_{\min},\\ \min\{|r_{i}-r_{j}|,T_{\max}\},&\text{if }T_{\min}\leq|r_{i}-r_{j}|.\end{cases}\end{split}$ (9)
@@
[…]
TF-IDF [SprckJones2021ASI] - 9.98 2.52 7.24 5.24 24.94
DSIR [Xie2023DataSF] - -0.02 0.49 2.01 -0.49 -2.10
[…]
@@ 4 Experimental Setup
Our experiments focus on LM finetuning, using the Qwen2.5 model family [Yang2024Qwen2TR] as our base LMs. During training, we start with the base LM and fine-tune it using a batch size of $32$ and the AdamW optimizer [Loshchilov2017DecoupledWD] with a learning rate of 2e-5 for two epochs.
@@
For AirRep training, we utilize two datasets, each representing distinct scenarios: instruction tuning on standard NLP tasks and training to improve conversational abilities: (i) FLAN [Wei2021FinetunedLM], an instruction-tuning dataset for language models, and (ii) UltraChat [Ding2023EnhancingCL], a large-scale dataset of instructional conversations covering a wide range of topics. To generate training signal, we set $N_{v}=10^{4}$ and $N_{t}=10^{5}$. The training subsets number is $M=100$, with each subsets containing n=1,000n=1{,}000 samples. We construct $100$ cross-validation instances. Thus, in total, the data includes $10^{4}$ unique training subsets and $10^{7}$ training examples. The Qwen2.5-0.5B LM is then fine-tuned on these training subsets and evaluated on the corresponding validation subset to obtain the label $r(\cdot,\cdot)$.
@@
We initialize AirRep using the GTE-Small, a 30M parameter embedding model [Li2023TowardsGT], and apply a randomly initialized projection matrix on top. AirRep is trained separately on FLAN and UltraChat. To construct the data for each training step of AirRep, we randomly select one cross-validation instance, then sample 1,0001{,}000 examples from its validation subset and $32$ training subsets. Distributed training is employed to maximize GPU memory utilization. The clipping thresholds, $T_{\text{min}}$ and $T_{\text{max}}$, are set to $0.1$ and $5.0$, respectively. The model is optimized for up to 2,0002{,}000 steps using the AdamW optimizer with a learning rate of $1\times 10^{-4}$.
@@
We use the following datasets for evaluation: (i) FLAN [Wei2021FinetunedLM]: The evaluation data of FLAN contains 66 NLP tasks spanning diverse categories. (ii) Alpaca [taori2023stanford]: An instruction-tuning dataset generated by OpenAI’s text-davinci-003 model. (iii) Tulu [Wang2023HowFC]: An instruction-tuning dataset comprising diverse data sources. (iv) SafeRLHF [Dai2023SafeRS]: A dataset for safety alignment of large language models, where each response is labeled as either safe or unsafe. Note that each evaluation data contain a test set and a training set. We ensure that all evaluation data remain excluded from AirRep’s optimization data to guarantee that the evaluation results reflect AirRep’s generalization to unseen data. Appendix D provides additional dataset details.
@@
We compare AirRep with representative gradient-based and representation-based methods. The gradient-based baselines we consider are as follows: (i) LoGra [Choe2024WhatIY], an optimized version of the influence function as described in Eq.4, based on its implementation in the Logix software library. (ii) TracIn [Pruthi2020EstimatingTD], which leverages the dot product of gradient vectors. (iii) LESS [Xia2024LESSSI], which computes projected gradients on LoRA weights and adjusts gradients using the AdamW states. (iv) Dsdm [Engstrom2024DsDmMD], an efficient implementation of Trak [Park2023TRAKAM] for LLM tasks.
The representation-based baselines we consider are as follows: (i) TF-IDF [SprckJones2021ASI], (ii) DSIR [Xie2023DataSF], hashed n-gram features, (iii) RDS [Hanawa2021EvaluationOS], which utilizes the last-layer hidden states of LMs, and (iv) GTE [Li2023TowardsGT], a text embedding model trained on text similarity and retrieval tasks which was also used as the base pre-trained model for AirRep. See Appendix E for more implementation details.